4. A 2.5 litre sample of a gas is heated from 250 C to 550C at constant pressure. Compute the volume at the higher temperature?




1
Expert's answer
2022-05-17T14:38:05-0400

Given:

V1 = 2.5 L

T1 = 25°C + 273.15 = 298.15 K

P = constant

V2 = unknown

T2 = 55°C + 273.15 = 328.15 K


Formula: V1/T1 = V2/T2


Solution:

Since the pressure and amount of gas remain unchanged, Charles's law can be used.

Charles's law can be expressed as: V1 / T1 = V2 / T2

V1T2 = V2T1

To find the volume at the higher temperature, solve the equation for V2:

V2 = V1T2 / T1

V2 = (2.5 L × 328.15 K) / (298.15 K) = 2.75 L

V2 = 2.75 L


Answer: 2.75 L